OpenStudy (anonymous):

A model rocket is launched with an initial upward velocity of 215 ft/s. The rocket's height h (in feet) after t seconds is given by the following. h=215t-16t^2 Find all values of t for which the rocket's height is 97 feet.

OpenStudy (amistre64):

h=215t-16t^2 97=215t-16t^2 0 = -97 + 215t - 16t^2 solve using your favorite quadratic method.

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

good luck